# Value-Penalized Auxiliary Control from Examples for Learning without Rewards or Demonstrations

### v1

@@ D.4 Additional Task Details
(a) Panda $\rho_{0}$ and $\mathcal{B}^{*}_{\mathcal{T}}$ examples.
(b) Sawyer and Adroit $\rho_{0}$ and $\mathcal{B}^{*}_{\mathcal{T}}$ examples.
(c) Real Panda $\rho_{0}$ and $\mathcal{B}^{*}_{\mathcal{T}}$ examples.
Figure 9: Examples of samples from initial state distribution $\rho_{0}$, auxiliary task buffers $\mathcal{B}^{*}_{\text{aux}}$, and main task buffers $\mathcal{B}^{*}_{\text{main}}$ for all environments and tasks used in this work. Auxiliary task data is reused between tasks in the Panda environment, since the observation space and action space are shared for all tasks, while only the task definitions themselves are reusable in the Panda and Adroit environments. Images are shown here for illustrative purposes, but all algorithms tested in this work use numerical state data.
D.5 Additional Algorithm Details
@@
22:        # Compute value penalization terms, see Section D.6.1
@@
VPACE-SQIL SQIL MSE Ours
ACE-SQIL SQIL MSE Ours
@@
Table 1: Major differences between algorithms studied in this work. MSE refers to mean squared error, while BCE refers to binary cross entropy.
Algorithm 1 shows a summary of VPACE, built on LfGP [4] and SAC-X [31]. Table 1 shows a breakdown of some of the major differences between all of the algorithms studied in this work.
D.6 Additional Implementation Details
@@
Total Interactions Task-specific (see Appendix E)
@@
$Q$ Update Freq. 1
Target $Q$ Update Freq. 1
$\pi$ Update Freq. 1
@@
SQIL Labels (Section D.6 $(-1,1)$
Expert Aug. Factor (Section D.6.3) 0.1
@@
$Q^{\pi}_{\text{max}}$, $Q^{\pi}_{\text{min}}$ num. filter points (Section D.6.1) 50
Auxiliary Control (ACE) Scheduler (Section D.6.2)
@@
Table 2: Hyperparameters shared between all algorithms, unless otherwise noted.
Whenever possible, all algorithms and baselines use the same modifications. Table 2 also shows our choices for common off-policy RL hyperparameters as well as choices for those introduced by this work.
DAC reward function: for VPACE-DAC and VP-DAC, although there are many options for reward functions that map $D$ to $\hat{R}$ [45], following [4, 5, 17], we set the reward to $\hat{R}_{\mathcal{T}}(s)=\log\left(D_{\mathcal{T}}(s)\right)-\log\left(1-D_{\mathcal{T}}(s)\right)$.
@@
Absorbing states and terminal states: for all algorithms, we do not include absorbing states (introduced in [5]) or terminal markers (sometimes referred to as “done”), since we found that both of these additions cause major bootstrapping problems when environments only terminate on timeouts, and timeouts do not necessarily indicate failure. Previous work supports bootstrapping on terminal states when they are caused by non-failure timeouts [36].
@@
D.6.1 Maintaining $Q^{\pi}_{\text{max}}$, $Q^{\pi}_{\text{min}}$ Estimates for Value Penalization
Our approach to value penalization requires maintaining estimates for or choosing $Q^{\pi}_{\text{max}}$ and $Q^{\pi}_{\text{min}}$. In both DAC and SQIL, the estimate of $Q^{\pi}_{\text{max}}$ comes from taking the mini-batch of data from $\mathcal{B}^{*}$, passing it through the Q function, taking the mean, and then using a median moving average filter to maintain an estimate. The “$Q^{\pi}_{\text{max}}$, $Q^{\pi}_{\text{min}}$ num. filter points” value from Table 2 refers to the size of this filter. We chose 50 and used it for all of our experiments. We set $Q^{\pi}_{\text{min}}$ to
@@ D.6.1 Maintaining $Q^{\pi}_{\text{max}}$, $Q^{\pi}_{\text{min}}$ Estimates for Value Penalization
D.6.2 Scheduler Choices
As noted in Section 4.1, our ACE algorithms use the same approach to scheduling from [4]. Specifically, we use a weighted random scheduler (WRS) combined a small set of handcrafted high-level trajectories. The WRS forms a prior categorical distribution over the set of tasks, with a higher probability mass $p_{\mathcal{T}_{\text{main}}}$ (Main Task Rate in Table 2) for the main task and $\frac{p_{\mathcal{T}_{\text{main}}}}{K}$ for all other tasks. Additionally, we choose whether to uniformly sample from a small set of handcrafted high-level trajectories, instead of from the WRS, at the Handcraft Rate from Table 2.
@@ D.6.2 Scheduler Choices
Figure 10: Additional results for the Unstack-Stack task from Fig. 11, either including or excluding the use of example augmentation described in Section D.6.3.
@@
D.6.3 Expert Data Augmentation
We added a method for augmenting our expert data to artificially increase dataset size. The approach is similar to other approaches that simply add Gaussian or uniform noise to data in the buffer [47, 48]. In our case, we go one step further than the approach from [48], and first calculate the per-dimension standard deviation of each observation in $\mathcal{B}^{*}$, scaling the Gaussian noise added to each dimension of each example based on the dimension’s standard deviation. For example, if a dimension in $\mathcal{B}^{*}$ has zero standard deviation (e.g., in Insert, the pose of the blue block is always the same), it will have no noise added by our augmentation approach. The parameter “Expert Aug. Factor” from Table 2 controls the magnitude of this noise, after our per-dimension normalization scheme.
@@ D.6.3 Expert Data Augmentation
D.6.4 Other Ablation Details
In our ablation experiments from Section 4.3, we included three baselines with full trajectory data, in addition to success examples. We added 200 $(s,a)$ pairs from full expert trajectories to make datasets comparable to the datasets from [4], where they used 800 expert $(s,a)$ pairs, but their environment was run at 20Hz instead of 5Hz, meaning they needed four times more data to have roughly the same total number of expert trajectories. We generated these trajectories using high-performing policies from our main experiments, since the raw trajectory data from [4] would not apply given that we changed the control rate from 20Hz to 5Hz.
D.6.5 Real Panda Implementation Details
While most of the design choices in Section D.6 apply to all environments tested, our real Panda environment had some small specific differences, mostly due to the complications of running reinforcement learning in the real world. We list the differences here, but for an exhaustive list, our open source code contains further details.
@@ D.6.5 Real Panda Implementation Details
Extra gradient steps: To add efficiency during execution and training, we completed training steps during the gap in time between an action being executed and an observation being gathered. Instead of completing a single gradient step at this time, as is the case for standard SAC (and VPACE), we completed four gradient steps, finding in simulated tasks that this gave a benefit to learning efficiency without harming performance. Previous work [4, 49] has found that increasing this update rate can reduce performance, but we hypothesize that our value penalization scheme helps mitigate this issue.
@@
Appendix E Additional Performance Results
@@
In this section, we expand upon our performance results from Section 4.2 and our Q-value calibration results from Section 4.4. We also show performance results for auxiliary tasks.
E.1 Expanded Main Task Performance Results
Fig. 11, Fig. 12, and Fig. 13 show expanded per-main-task results for our Panda (originally presented in [4, 32]), Sawyer (originally presented in [11, 3]), and Adroit Hand (originally presented in [37, 3]) tasks, respectively.
Figure 11: Performance results on simulated Panda tasks from [4, 32]. Performance on auxiliary tasks is not shown. Methods that leverage the multitask framework are shown with solid lines, and their single-task counterparts are shown with dashed lines. The shaded area corresponds to half standard deviation across five seeds. Both value-penalization and auxiliary task exploration are consistently beneficial for the most complicated tasks.
Figure 12: Performance results on the Sawyer tasks considered in [3] and introduced in [11]. For more complex tasks, the multitask methods learn with far fewer environment steps. These results are more stochastic than those from Fig. 11, but the same overall pattern emerges.
Figure 13: Performance results on the Adroit tasks considered in [3] and introduced in [37], along with modified versions of the environments that use delta-position based control. The episode return doesn’t fully represent the results for these environments, because the policies learned for the original tasks can be quite coarse and unrealistic (see Fig. 7). For the original environments, many baselines perform comparably, but for the dp versions, VPACE-SQIL, VP-SQIL, and VPACE-DAC are the most consistently high performing.
In the Panda tasks (Fig. 11), the hardest tasks (Stack, Unstack-Stack, and Insert) benefit the most from VPACE, but VPACE-SQIL is always the fastest and highest performing method. Notably, ACE-SQIL is always outperformed by both VPACE-SQIL and VPACE-DAC, and RCE is always outperformed by VP-SQIL. The benefits of value penalization are also clear, even in Bring, where ACE-SQIL intially learns the task, but starts to learn very poor Q estimates and have increasingly poorer and more unstable performance.
In the Sawyer tasks (Fig. 12), the benefit of VPACE is not as prominent, even if VPACE-SQIL is still the best performing method on average. The clearer pattern is the improvement from adding auxiliary task exploration, although the improvement is still lower, on average, than the improvement in the Panda tasks. Compared with the Panda tasks, these tasks do not have randomized initial conditions (apart from sawyer_box_close), and also generally require shorter movements to complete, so the benefits of VPACE are less pronounced.
In the Adroit hand tasks (Fig. 13), VPACE-SQIL, VPACE-DAC, and VP-SQIL are always among the highest performing methods. In particular, RCE and ACE-RCE are always outperformed by all other methods. As stated in Section 4.1 and further detailed in Section D.2, we also generated delta-position variants of the environments, which (i) generated, qualitatively, far more realistic policies (see Fig. 7), and (ii) allowed many methods to solve the relocate task.
E.2 Value Penalization Calibration Improvement – All Environments
Figure 14: Additional results, from the Panda (left) and Sawyer (right) tasks, showing the difference between per-timestep Q values and the average Q values for data from $\mathcal{B}^{*}_{\text{main}}$. The most difficult tasks, with a larger variety of possible states, tend to have violations of the ideal maximum when value penalization is excluded, often exacerbated when ACE is used.
@@ E.2 Value Penalization Calibration Improvement – All Environments
This section expands on the value penalization calibration results from Section 4.4 for the Panda and Sawyer tasks Fig. 14 and for the Adroit hand tasks Fig. 15. Following the results shown for Unstack-Stack, the results for other difficult tasks also show clear violations of $y(s,a)\leq y(s^{*},a^{*})$, and tasks in which this rule is violated also tend to have poorer performance. The violations are more pronounced ACE-SQIL, which, as previously discussed, is likely because they generate far more diverse state distributions in $\mathcal{B}$. As shown for Unstack-Stack in Section 4.4, VP-SQIL and VPACE-SQIL never violate $y(s,a)\leq y(s^{*},a^{*})$. Intriguingly, although the rule is severely violated for many Adroit hand tasks, ACE-SQIL and SQIL still have reasonable performance in some cases. This shows that highly uncalibrated Q estimates can still, sometimes, lead to adequate performance. We hypothesize that this occurs because these tasks do not necessarily need $s_{T}\sim\mathcal{B}$ to match $s^{*}\sim\mathcal{B}^{*}_{\text{main}}$ to achieve high return, but we leave investigating this point to future work.
E.3 Auxiliary Task Performance Results
Figure 15: Additional results from the Adroit tasks showing the difference between per-timestep Q values and the average Q values for data from $\mathcal{B}^{*}_{\text{main}}$.
Figure 16: Auxiliary task performance for all of our methods that employ auxiliary control from examples. Main tasks are divided by row, and auxiliary tasks are divided by column, with the right-most column in each row showing main task performance. Performance on auxiliary tasks mostly follows performance on the main task, indicating that these auxiliary tasks may be useful for transfer learning to other main tasks.
Figure 17: Sawyer auxiliary task performance. In the Sawyer tasks, auxiliary task varies but is often reasonable for methods with value penalization, and poorer when it is excluded.
Figure 18: Adroit hand auxiliary task performance. Particularly in the hand environments, auxiliary task performance can be quite poor, which we expect is because we evaluate auxiliary tasks under different circumstances from how they are used during learning.
In our main text, we only presented performance results for main tasks, since our primary goal was to efficiently generate policies that performed well on a particular task. However, our scheduled auxiliary task approach inherently learns multiple auxiliary task policies $\pi_{\mathcal{T}_{\text{aux}}}$ in addition to the main task policy $\pi_{\mathcal{T}_{\text{main}}}$. While performance on these auxiliary task policies is not necessarily expected to be high throughout learning, observing auxiliary task performance can potentially provide useful insights, such as whether the auxiliary policies would be useful for transfer learning to learn a new main task.
We show all auxiliary task performance in for our Panda environments in Fig. 16. Auxiliary tasks tend to be learned slightly before the main task is learned, and otherwise auxiliary task performance generally follows performance on the main task (i.e. if the main task is learned slowly or poorly, auxiliary task performance will reflect this). The performance on the Reach auxiliary task is sometimes lower than on other, ostensibly more difficult tasks such as Lift, but this is likely because the success criteria for Reach is less forgiving than that for Lift.
We show auxiliary task performance for Sawyer and Adroit Hand environments in Fig. 17 and Fig. 18. Returns for auxiliary tasks are calculated based on reward functions originally taken from [3], in which a reaching reward is positive as an end-effector reaches closer to an object, and negative as it moves away, and grasp receives a similar bonus for moving or lifting the object of interest. In our ACE experiments for Sawyer and Adroit, we only use auxiliary tasks as part of a handcrafted high-level trajectory (see Section D.6.2), meaning that reach and grasp are only learned after a reset, whereas during evaluation, the policy may have an opportunity to guide the end-effector to partially reach an object, then move away, with no learned strategy for recovery. We suspect that using a weighted random scheduler throughout learning may alleviate this issue, but we did not explore this in this work, since learning generally effective or reusable auxiliary tasks is beyond the scope of this work.
Appendix F Expanded Limitations
In this section, we expand on some of the limitations originally discussed in Section 5, where we previously skipped limitations that are inherent to reinforcement learning and to learning from guided play (LfGP, [4]), both of which are core parts of our approach.
Experimental limitation—Numerical state data only All of our tests are done with numerical data, instead of image-based data. Other work [31, 47] has shown that for some environments, image-based learning just results in slowing learning compared with numerical state data, and we assume that the same would be true for our method as well.

### v2

@@ B-D Additional Task Details
B-E Additional Algorithm Details
@@
22:        # Compute value penalization terms, see Section B-F1
@@
VPACE SQIL MSE Ours
ACE SQIL MSE Ours
@@
TABLE III: Major differences between algorithms studied in this work. MSE refers to mean squared error, while BCE refers to binary cross entropy.
Algorithm 1 shows a summary of VPACE, built on LfGP [6] and SAC-X [5]. As a reminder, our multi-policy objective function is
$\mathcal{L}(\pi;\mathcal{T})=\mathbb{E}_{\mathcal{B},\pi_{\mathcal{T}}}\left[Q_{\mathcal{T}}(s,a)\right],$ (11)
our multi-q objective function is
$\begin{split}\mathcal{L}(Q;\mathcal{T})=&\mathbb{E}_{\mathcal{B},\pi_{\mathcal{T}}}\left[(Q_{\mathcal{T}}(s,a)-y_{\mathcal{T}}(s,s^{\prime}))^{2}\right]\\ &+\mathbb{E}_{\mathcal{B}^{*}_{\mathcal{T}},\pi_{\mathcal{T}}}\left[(Q_{\mathcal{T}}(s^{*},a)-y_{\mathcal{T}}(s^{*},s^{*}))^{2}\right],\end{split}$ (12)
and our value-penalized update is
$\begin{split}\mathcal{L}^{\pi}_{\text{pen}}(Q)=\lambda\mathbb{E}_{\mathcal{B}}[&\left(\max(Q(s,a)-Q^{\pi}_{\text{max}},0)\right)^{2}\\ &+\left(\max(Q^{\pi}_{\text{min}}-Q(s,a),0)\right)^{2}],\end{split}$ (13)
where
$Q^{\pi}_{\text{min}}=\hat{R}_{\text{min}}/(1-\gamma),$ (14)
and
$Q^{\pi}_{\text{max}}=\mathbb{E}_{\mathcal{B}^{*}}\left[V^{\pi}(s^{*})\right].$ (15)
Table III shows a breakdown of some of the major differences between all of the algorithms studied in this work.
B-F Additional Implementation Details
@@
TABLE IV: Hyperparameters shared between all algorithms, unless otherwise noted.
@@
Total Interactions Task-specific (see Appendix C)
@@
$Q$ Update Freq. 1 (sim), 4 (real)
Target $Q$ Update Freq. 1 (sim), 4 (real)
$\pi$ Update Freq. 1 (sim), 4 (real)
@@
SQIL Labels (Section B-F $(-1,1)$
Expert Aug. Factor (Section B-F3) 0.1
@@
$Q^{\pi}_{\text{max}}$, $Q^{\pi}_{\text{min}}$ num. filter points (Section B-F1) 50
Auxiliary Control (ACE) Scheduler (Section B-F2)
@@
Whenever possible, all algorithms and baselines use the same modifications. Table IV also shows our choices for common off-policy RL hyperparameters as well as choices for those introduced by this work.
DAC reward function: for VPACE-DAC and VP-DAC, although there are many options for reward functions that map $D$ to $\hat{R}$citeappghasemipourDivergenceMinimizationPerspective2019, following [6, 14, 17], we set the reward to $\hat{R}_{\mathcal{T}}(s)=\log\left(D_{\mathcal{T}}(s)\right)-\log\left(1-D_{\mathcal{T}}(s)\right)$.
@@
[…]
@@
B-F1 Maintaining $Q^{\pi}_{\text{max}}$, $Q^{\pi}_{\text{min}}$ Estimates for Value Penalization
Our approach to value penalization requires maintaining estimates for or choosing $Q^{\pi}_{\text{max}}$ and $Q^{\pi}_{\text{min}}$. In both DAC and SQIL, the estimate of $Q^{\pi}_{\text{max}}$ comes from taking the mini-batch of data from $\mathcal{B}^{*}$, passing it through the Q function, taking the mean, and then using a median moving average filter to maintain an estimate. The “$Q^{\pi}_{\text{max}}$, $Q^{\pi}_{\text{min}}$ num. filter points” value from Table IV refers to the size of this filter. We chose 50 and used it for all of our experiments. We set $Q^{\pi}_{\text{min}}$ to
@@ B-F1 Maintaining $Q^{\pi}_{\text{max}}$, $Q^{\pi}_{\text{min}}$ Estimates for Value Penalization
B-F2 Scheduler Choices
Our ACE algorithms use the same approach to scheduling from [6]. Specifically, we use a weighted random scheduler (WRS) combined a small set of handcrafted high-level trajectories. The WRS forms a prior categorical distribution over the set of tasks, with a higher probability mass $p_{\mathcal{T}_{\text{main}}}$ (Main Task Rate in Table IV) for the main task and $\frac{p_{\mathcal{T}_{\text{main}}}}{K}$ for all other tasks. Additionally, we choose whether to uniformly sample from a small set of handcrafted high-level trajectories, instead of from the WRS, at the Handcraft Rate from Table IV.
@@ B-F2 Scheduler Choices
Figure 14: Variations of example augmentation. Our example augmentation scheme provides a small but noticeable bump in performance, which is magnified with when fewer expert examples are used.
@@
B-F3 Expert Data Augmentation
[…]
@@ B-F3 Expert Data Augmentation
B-F4 Other Ablation Details
In our ablation experiments from, we included three baselines with full trajectory data, in addition to success examples. We added 200 $(s,a)$ pairs from full expert trajectories to make datasets comparable to the datasets from [6], where they used 800 expert $(s,a)$ pairs, but their environment was run at 20Hz instead of 5Hz, meaning they needed four times more data to have roughly the same total number of expert trajectories. We generated these trajectories using high-performing policies from our main experiments, since the raw trajectory data from [6] would not apply given that we changed the control rate from 20Hz to 5Hz.
B-F5 Real Panda Implementation Details
Figure 15: Sample efficiency performance plots for main task only for all simulated main tasks, including baselines not shown in the main text. Methods introduced in this work have solid lines, while baselines are shown with dashed lines. Performance is an interquartile mean (IQM) across 5 timesteps and 5 seeds with shaded regions showing 95% stratified bootstrap confidence intervals [7].
Figure 16: Comparison of underlying reward models when used with ACE. Our main results (VPACE) used SQIL as the reward model. VPACE-DAC eventually reaches similar performance to VPACE-SQIL, but tends to take longer, and ACE-RCE often has far poorer performance than both.
Figure 17: Comparison of different variants of SQIL: both VP and ACE (VPACE), VP-only (VP-SQIL), or SQIL alone. VP-SQIL generally either results in an improvement or has no effect on performance, compared with SQIL, but the exclusion of ACE results in far poorer performance than VPACE, especially for the most complex tasks.
Figure 18: Comparison of different variants of SQIL: both VP and ACE (VPACE), VP-only (VP-SQIL), or SQIL alone. VP-SQIL generally either results in an improvement or has no effect on performance, compared with SQIL, but the exclusion of ACE results in far poorer performance than VPACE, especially for the most complex tasks.
While most of the design choices in Section B-F apply to all environments tested, our real Panda environment had some small specific differences, mostly due to the complications of running reinforcement learning in the real world. We list the differences here, but for an exhaustive list, our open source code contains further details.
@@ B-F5 Real Panda Implementation Details
[…]
@@
Appendix C Additional Performance Results
@@
In this section, we expand upon our performance results and our Q-value overestimation analysis.
C-A Expanded Main Task Performance Results
Fig. 15 shows expanded results for our simulated environments, with each baseline shown for each individual environment.
C-B ACE Reward Model Comparison
The SAC-X framework, which we describe as ACE when used with example states only, is agnostic to the choice of reward model (see Appendix A). Therefore, we also completed experiments in each of our simulated environments with DAC and with RCE as the base reward model, instead of SQIL, which was used for all of our main VPACE results. The results are shown in Fig. 16, where it is clear that VPACE with SQIL learns more efficiently than VPACE with DAC and with higher final performance than ACE with RCE. We do not use value penalization (VP) for RCE, since RCE uses a classification loss for training Q that would not benefit from the use of value penalization.
C-C Single-task Reward Model Comparison
Figure 19: Results for different reward models for EBC, excluding the use of both value penalization (VP) and auxiliary control from examples (ACE).
[…]
C-D Single-task Value Penalization
Our scheme for value penalization, while initially motivated by the use of ACE, can be used in the single task regime. In Fig. 17, we compare the performance of VPACE, SQIL, and SQIL with value penalization, but without auxiliary tasks (VP-SQIL). VP-SQIL either has no effect on performance or results in an improvement over SQIL, as expected, but is still strongly outperformed by VPACE on the more complicated tasks, such as Stack, Unstack-Stack, Insert, sawyer_box_close, sawyer_bin_picking, and the dp variant of relocate-human-v0.
C-E Q-Value Overestimation and Penalization – All Environments
@@
Fig. 18 shows the same value overestimation analysis plots shown in the original paper for all tasks in. The results for other difficult tasks also show clear violations of $y(s,a)\leq y(s^{*},a^{*})$, and tasks in which this rule is violated also often have poorer performance. Intriguingly, although the rule is severely violated for many Adroit hand tasks, ACE without VP still has reasonable performance in some cases. This shows that highly uncalibrated Q estimates can still, sometimes, lead to adequate performance. We hypothesize that this occurs because these tasks do not necessarily need $s_{T}\sim\mathcal{B}$ to match $s^{*}\sim\mathcal{B}^{*}_{\text{main}}$ to achieve high return, but we leave investigating this point to future work.
Appendix D Why Does SQIL Outperform RCE?
[3] claimed that the highest performing baseline against RCE was SQIL, but it is worth noting that their implementation [^1] Available at https://github.com/google-research/google-research/tree/master/rce at time of writing. is a departure from the original SQIL implementation, which uses MSE for TD updates, described in Section A-B. Furthermore, [3] also noted that it was necessary to add $n$-step returns to off-policy learning to get reasonable performance. In their experiments, however, this adjustment was not included for their version of SQIL with BCE loss. We complete the experiments using their implementation and show the average results across all RCE environments in Fig. 20, and also compare to using SQIL with MSE (without value penalization or auxiliary tasks) for TD updates.
These results empirically show that RCE and SQIL with BCE loss perform nearly identically, indicating that benefits of the changed TD targets and weights described in Section A-C may not be as clear as previously described. SQIL with MSE clearly performs better on average, although it still performs worse than VPACE (see Appendix C).
Appendix E Limitations
In this section, we detail limitations of this work.
[…]